OpenStudy (anonymous):

Solve 3 log6 2 + 2 log6 3 – log6 x = 2

OpenStudy (anonymous):

little 6's

OpenStudy (anonymous):

do you know the log properties??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so using that i am going to help you find x....

OpenStudy (anonymous):

\[a \log _{b}x= \log _{b}x ^{a}\]

OpenStudy (anonymous):

using that for your equation, what do you get??

OpenStudy (anonymous):

idk i dont get it with constants

OpenStudy (anonymous):

try it again.....in the case of the constants, if you pick \[3 \log _{6}2\] your a is 3 , your b is 6 , and your 2 is your x....

OpenStudy (anonymous):

do you get it??

OpenStudy (anonymous):

no

OpenStudy (anonymous):